A person who weighs 644 N is riding a 87-N mountain bike. Suppose the entire weight of the rider plus bike is supported equally by the two tires. If the gauge pressure in each tire is 5.60 x 105 Pa, what is the area of contact between each tire and the ground?
Answer #1

F = ( P ) ( A )

A = F / P

A = ( 644+ 87 ) / (560,000) = 1.3x 10^-3 sq. m. <------------------------------